OpenStudy (anonymous):

Given points A(0,2) and B (-5,15) when a right triangle is drawn to find the distance between A or B, Find the length of a horizontal leg using Pythagorean thereom. If can , can some draw to scale so i could grasp it better

OpenStudy (anonymous):

@Calcmathlete hi. could you draw this one like you drew the other . the only way i could grasp this

OpenStudy (campbell_st):

this question makes no sense.... you don't need pythagoras's theorem to find the horizontal leg...

OpenStudy (anonymous):

You don't really need the drawing to do this kind of thing. JUst find the distance between the x-values and y-values to get the two legs of the pythagorean theorem. |0 - - 5| = 5 |2 - 15| = 13 The two legs are 5 and 13. Here's a drawing anyway :) 

OpenStudy (anonymous):

Actually, yeah. I agree with @campbell_st I solved the horizontal leg without Pythagorean Theorem. TO use it to solve it involves a lot more work...

OpenStudy (anonymous):

sooo how do i do it , if not PT ?

OpenStudy (campbell_st):

you can find the hypotenuse ( longest side) of the triangle using pythagoras' theorem... thats about it

OpenStudy (anonymous):

Take a look at the part above my drawing. Subtract the x-values and find the absolute value of that. |0 - - 5| = = |0 + 5| = 5

OpenStudy (anonymous):

That's how I'd do it without Pythagorean Theorem...it's more like a preparation for Pythagorean THeorem...

OpenStudy (anonymous):

my answer was 13 . am i right

OpenStudy (anonymous):

Again, I just showed it to you right here: |0 - - 5| = = |0 + 5| = 5

OpenStudy (campbell_st):

nope... the answer for the length AB is \[AB = \sqrt{194}\]